# Communication Cost Reduction for Subgraph Counting under Local Differential Privacy via Hash Functions

### v1

@@ Abstract
We suggest the use of hash functions to cut down the communication costs when counting subgraphs under edge local differential privacy. While various algorithms exist for computing graph statistics, including the count of subgraphs, under the edge local differential privacy, many suffer with high communication costs, making them less efficient for large graphs. Though data compression is a typical approach in differential privacy, its application in local differential privacy requires a form of compression that every node can reproduce. In our study, we introduce linear congruence hashing. With a sampling rate of $s$, our method can cut communication costs by a factor of $s^{2}$, albeit at the cost of increasing variance in the published graph statistic by a factor of $s$. The experimental results indicate that, when matched for communication costs, our method achieves a reduction in the $\ell_{2}$-error for triangle counts by up to 1000 times compared to the performance of leading algorithms.
I Introduction
@@
Differential privacy [1, 2] has emerged as a benchmark for protecting user data. To meet this standard, it is necessary to obfuscate the publication results. This can be achieved by adding small noise [3] or altering the publication outcomes with small probability [4].
@@
In differential privacy, it is typically assumed that there is a complete dataset available. After statistical analyses are done, obfuscation is applied to the results. However, there can be leaks of user information during the data collection or storage phases. To address these concerns, a variant of differential privacy, termed local differential privacy [5, 6], has been introduced. Here, rather than starting with a full dataset, each user is prompted to disguise their data before sharing it. As a result, the received dataset is not perfect. Numerous studies, like those referenced in [7, 8], have aimed to extract accurate statistical insights from these imperfect datasets. The local differential privacy has been practically employed by several companies to guarantee the privacy of their users’ information. Those companies include Apple, Microsoft, and Google [9, 10, 11].
@@
Many studies on local differential privacy focus on tabular datasets, but there is also significant research dedicated to publishing graph statistics [12, 13]. For graph-based inputs, such as social networks, the prevalent privacy standard is edge local differential privacy [14]. In this framework, users are asked to share a disguised version of their adjacency vectors. These vectors, essentially bit sequences, signify mutual connections within the network. For instance, in a social network comprising $n$ users, user $v_{i}$ would provide their adjacency vector, denoted as $a_{i}=[a_{i,1},\dots,a_{i,n}]\in\{0,1\}^{n}$. Here, $a_{i,j}=1$ implies that users $v_{i}$ and $v_{j}$ are connected, while $a_{i,j}=0$ means there is no connection between them.
@@
One widely adopted method for obfuscation is the randomized response [15, 16, 17]. In this approach, users are prompted to invert each bit in their adjacency vector, denoted as $a_{i,j}$, based on a specific probability. However, while the method is straightforward, its application to real-world social networks presents challenges. Typically, in practical social networks, an individual may have at most a few thousand friends, implying that the majority of $a_{i,j}$ values are zeroes. When employing the randomized response technique and flipping each bit of $a_{i,j}$ based on the designated probability, a significant number of zero bits get inverted to one. This phenomenon can distort the resulting graph statistics considerably [18, 19].
@@
Beyond the publication of full graphs via randomized rounding, there is a growing interest in subgraph counting queries under local differential privacy. The challenge was first highlighted by [20], who proposed two key algorithms for $k$-star and triangle counting. Building on the $k$-star counting, [21] later introduced an unbiased and refined algorithm. For triangle counting, several noteworthy contributions include [22], which addressed the communication cost, [23], which delved into scenarios where users have access to a 2-hop graph view, and [24], which studies the lower bound of the additive error. Moving to the shuffle model, [25] gave algorithms for both triangle and 4-cycle counting. To conclude, [26] gave an approximation method for identifying the densest subgraph, rooted in locally adjustable algorithms.
@@
Most of the works mentioned in the previous paragraph use the two-step mechanism [20, 22, 25]. In this mechanism, users are required to download the adjacency vectors of all other participants. They then compute graph statistics locally using their genuine adjacency vector and the obfuscated vectors from their peers. While this technique substantially boosts the accuracy for various graph statistics, such as the count of triangles, it comes with the drawback of demanding a vast amount of data download. Each user incurs a communication cost of $\Theta(n^{2})$. Given that the user count, $n$, can reach several billion in real-world scenarios, this download demand is not affordable for the majority.
@@
Instead of using adjacency vectors, users might consider exchanging adjacency lists. This approach involves each user transmitting and receiving only a list of node pairs that are adjacent. If we assume that a user can have a limited, constant number of friends, then the bit-download requirement for the original graph becomes $\Theta(n\log n)$. However, given the surge in edge count to $\Theta(n^{2})$ due to the randomized response, the communication cost associated with downloading the obfuscated adjacency list shoots up to $\Theta(n^{2}\log n)$. This means that, post-obfuscation, the adjacency list fails to offer any communication efficiency benefits over the adjacency vector.
@@
There are several techniques proposed to reduce the communication cost for the two-step mechanism such as the asymmetric randomized response [22], the degree-preserving randomized response [27], the degree-preserving exponential mechanism [28], or technique based on the compressive sensing [29]. However, we strongly believe that the communication cost could be further reduced both in theory and practice.
I-A Our Contributions
@@
The concept of database compression is not novel within the realm of differential privacy. Numerous mechanisms have been developed to enhance the precision of statistics derived from tabular data through database compression using sampling. One of the most prominent among these is the SmallDB algorithm [30]. There have also been initiatives that harness the Kronecker graph to refine the accuracy of published graph statistics [31].
On the contrary, implementing database compression for the two-step mechanism under edge local differential privacy poses greater challenges. Imagine we compress the adjacency list $[a_{i,1},\dots,a_{i,n}]^{t}$ down to $[a_{i,c_{1}},\dots,a_{i,c_{m}}]^{t}$, with $m\ll n$, via sampling, and then disseminate this compressed list to all users. When other users utilize this compressed list for local calculations in the second step, they must be able to replicate the sampling outcome. We therefore need a function which can be replicated with small communication cost. This requirement inspired our proposition to utilize linear congruence hashing [32, 33] for the sampling process.
Using linear congruence hashing, we can evenly partition the node set $\{v_{1},\dots,v_{n}\}$ into $m$ subsets. Let us denote these subsets as $S_{1},\dots,S_{m}$, each having a size of $s$. The index $c_{i}$ is derived from $S_{i}$ with a uniform probability of $1/s$. Since all data is transmitted from the users with a probability of $1/s$, we can leverage the theorem on amplification by sub-sampling [34] to diminish the bit-flipping probability in the randomized response mechanism by a factor of $1/s$. Consequently, the count of non-zero entries in the bit vector $[a_{i,c_{1}},\dots,a_{i,c_{m}}]^{t}$ is roughly $m/s$, which is approximately equal to $n/s^{2}$. By employing a sampling rate of $s$, the communication overhead can be reduced by a factor of $s^{2}$.
@@
I-B Paper Organization
In Section II, we establish the fundamental notations of edge differential privacy and present the primitive mechanism that adheres to this privacy standard. Moving to Section III, we delve into the two-step mechanism, highlighting its communication costs in detail. Our approach, termed the group randomized response, is introduced in Section IV, which is then succeeded by a theoretical analysis in Section V. The methodology is applied to triangle counting in Section VI. We summarize and conclude our findings in Section VII.
II Preliminaries
II-A Hash functions
[…]
Each value in the hash space is referred to as a bin. Two keys, $\mathsf{x}$ and $\mathsf{y}$, are said to be in the same bin or to ‘collide’ if they have identical hash values. In other words, for a hash function $h$, a collision occurs if $h(\mathsf{x})=h(\mathsf{y})$. Hash functions are crafted to minimize the likelihood of collisions. Universal hashing, detailed in the subsequent definition, is a primary method employed in the design of these functions.
Definition 1 (Universal Hashing Scheme [36]).
A family of hash functions $\mathcal{H}$ is universal if, for all $\mathsf{x},\mathsf{y}\in[|0,n-1|]$ such that $\mathsf{x}\neq\mathsf{y}$, we have |$|\{h\in\mathcal{H}|h(\mathsf{x})=h(\mathsf{y})\}|\leq\frac{|\mathcal{H}|}{m}$.
@@ II-A Hash functions
It is shown in [36] that a scheme called linear congruence hashing [32, 33] is universal. The scheme can be defined as =$\mathcal{H}=\{h_{\mathsf{a},\mathsf{b}}|\mathsf{a}\in[|1,\mathsf{p}-1|],\mathsf{b}\in[|0,\mathsf{p}-1|]\}$ when $\mathsf{p}$ is a prime number greater than $n$ and $h_{\mathsf{a},\mathsf{b}}(\mathsf{x})=((\mathsf{a}\mathsf{x}+\mathsf{b})\text{mod }\mathsf{p})\text{ mod }m$. Notice that, for all $\mathsf{a,b}$ and $\mathsf{k}\in[|0,m-1|]$, we have $\left\lfloor\frac{n}{m}\right\rfloor\leq|\{\mathsf{x}:h_{\mathsf{a,b}}(\mathsf{x})=k\}|\leq\left\lceil\frac{n}{m}\right\rceil$. In other words, the number of key in a given bin is either $\left\lfloor\frac{n}{m}\right\rfloor$ or $\left\lceil\frac{n}{m}\right\rceil$.
II-B Edge Local Differential Privacy
@@
Let the set of users be $\{v_{1},\dots,v_{n}\}$. Each user $v_{i}$ has own adjacency list $a_{i}=[a_{i,1},\dots,a_{i,n}]^{t}\in\{0,1\}^{n}$ where $a_{i,j}=1$ if $\{i,j\}\in E$ and $a_{i,j}=0$ otherwise. For any $a_{i},a_{i}^{\prime}\in\{0,1\}^{n}$, let $d(a_{i},a_{i}^{\prime})$ be the $\ell_{1}$-distance between the two vectors. The edge local differential privacy can be explained as in the following definition:
Definition 2 (Edge Local Differential Privacy [14]).
Let $\varepsilon>0$, let :$\mathcal{R}:\{0,1\}^{n}\rightarrow R$ be a randomized function, and let $\mathcal{A}:R^{n}\rightarrow\mathcal{S}$ be a function called aggregator. The mechanism $\mathcal{M}$ which outputs $\mathcal{A}(\mathcal{R}(a_{1}),\ldots,\mathcal{R}(a_{n}))$ is said to be $\varepsilon$-edge local differentially private if, for any $a_{i},a_{i}^{\prime}\in\{0,1\}^{n}$ such that $d(a_{i},a_{i}^{\prime})=1$ and any $\mathsf{R}\subset R$, we have $\mathbb{P}\left(\mathcal{R}(a_{i})\in\mathsf{R}\right)\leq e^{\varepsilon}\mathbb{P}\left(\mathcal{R}(a_{i}^{\prime})\in\mathsf{R}\right)$.
@@
Theorem 1 (Composition Theorem [37]).
@@
II-C Basic Mechanisms
Numerous mechanisms have been suggested to meet Definition 2 [38, 39]. Among them, the edge local Laplacian mechanism [21] is an approach to offer privacy when we request each user to give real numbers to the aggregator. The Laplacian mechanism can be defined as in the following definition. It can be shown that the mechanism is $\varepsilon$-edge local differential private.
Definition 3 (Edge Local Laplacian Mechanism [21]).
For $f:\{0,1\}^{n}\rightarrow\mathbb{R}$, the global sensitivity of $f$ is $\Delta_{f}=\max\limits_{d(a,a^{\prime})=1}\lVert f(a)-f(a^{\prime})\rVert_{1}$. For $\varepsilon>0$, let $X$ be a random variable drawn from $\text{Lap}\left(\varepsilon/\Delta_{f}\right)$. The edge local Laplacian mechanism for $f$ with privacy budget $\varepsilon$ is the mechanism for which the randomized function $\mathcal{R}$ is defined as $f(a)+X$.
When we request users to send a vector of bits (such as the whole adjacency vector) to the aggregator, the edge local Laplacian mechanism is not the most suitable. Instead, a frequently used method is the randomized response [15, 16, 17], as detailed in the following definition.
Definition 4 (Randomized Response [15, 16, 17]).
For $\varepsilon>0$, the randomized response with privacy budget $\varepsilon$ is a mechanism which the randomized function :$\mathcal{R}:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ is a function such that $\mathcal{R}(a_{i,1},\dots,a_{i,n})=(RR(a_{i,1}),\dots,RR(a_{i,n}))$ where
$\mathbb{P}\left(\text{RR}(a_{i,j})=1\right)=\begin{cases}\frac{e^{\varepsilon}}{1+e^{\varepsilon}}&\text{if }a_{i,j}=1\\ \frac{1}{1+e^{\varepsilon}}&\text{if }a_{i,j}=0.\end{cases}$ (1)
II-D Smooth Sensitivity
The traditional Laplacian Mechanism adjusts noise based on the global sensitivity, which is determined from the worst possible outcome for a selected function. It does not consider the specific adjacency list. Therefore, in certain situations, the global sensitivity might be much greater than the function’s changes due to slight update in the real adjacency list. Noticing this, [40] created a mechanism where the noise relates to the smooth sensitivity, which varies depending on the instance. But before we define smooth sensitivity, we first need to explain local sensitivity and sensitivity at a distance $k$.
Definition 5 (Local Sensitivity from Definition 1.4 of [40]).

### v2

@@ Abstract
We suggest the use of hash functions to cut down the communication costs when counting triangle and other subgraphs under edge local differential privacy. While various algorithms exist for computing graph statistics — including the count of subgraphs — under the edge local differential privacy, many suffer with high communication costs, making them less efficient for large graphs. Though data compression is a typical approach in differential privacy, its application in local differential privacy requires a form of compression that every node can reproduce. In our study, we introduce linear congruence hashing. Leveraging amplification by sub-sampling, with a sampling size of $s$, our method can cut communication costs by a factor of $s^{2}$, albeit at the cost of increasing variance in the published graph statistic by a factor of $s$. The experimental results indicate that, when matched for communication costs, our method achieves a reduction in the $\ell_{2}$-error by up to 1000 times for triangle counts and by up to $10^{3}$ times for 4-cycles counts compared to the performance of leading algorithms.
1 Introduction
@@
Differential privacy (Dwork, 2006; Dwork et al., 2014) has emerged as a benchmark for protecting user data. To meet this standard, it is necessary to obfuscate the publication results. This can be achieved by adding small noise (Dwork et al., 2006) or altering the publication outcomes with small probability (McSherry & Talwar, 2007).
@@
In differential privacy, it is typically assumed that there is a complete dataset available. After statistical analyses are done, obfuscation is applied to the results. However, there can be leaks of user information during the data collection or storage phases. To address these concerns, a variant of differential privacy, termed local differential privacy (Cormode et al., 2018; Evfimievski et al., 2003), has been introduced. Here, rather than starting with a full dataset, each user is prompted to disguise their data before sharing it. As a result, the received dataset is not perfect. Numerous studies, like those referenced in Li et al. (2020); Asi et al. (2022), have aimed to extract accurate statistical insights from these imperfect datasets. The local differential privacy has been practically employed by several companies to guarantee the privacy of their users’ information. Those companies include Apple, Microsoft, and Google (Differential Privacy Team, Apple, 2017; Ding et al., 2017; Erlingsson et al., 2014).
@@
Many studies on local differential privacy focus on tabular datasets, but there is also significant research dedicated to publishing graph statistics (Sajadmanesh & Gatica-Perez, 2021; Ye et al., 2020). For graph-based inputs, such as social networks, the prevalent privacy standard is edge local differential privacy (Qin et al., 2017). In this framework, users are asked to share a disguised version of their adjacency vectors. These vectors are bit sequences which show mutual connections within the network. For instance, in a social network comprising $n$ users, user $v_{i}$ would provide their adjacency vector, denoted as $a_{i}=[a_{i,1},\dots,a_{i,n}]\in\{0,1\}^{n}$. Here, $a_{i,j}=1$ implies that users $v_{i}$ and $v_{j}$ are connected, while $a_{i,j}=0$ means there is no connection between them.
@@
One widely adopted method for obfuscation is the randomized response (Warner, 1965; Mangat, 1994; Wang et al., 2016). In this approach, users are prompted to invert each bit in their adjacency vector, denoted as $a_{i,j}$, based on a specific probability. However, while the method is straightforward, its application to real-world social networks presents challenges. Typically, in practical social networks, an individual may have at most a few thousand friends, implying that the majority of $a_{i,j}$ values are zeroes. When employing the randomized response technique and flipping each bit of $a_{i,j}$ based on the designated probability, a significant number of zero bits get inverted to one. This phenomenon can distort the resulting graph statistics considerably (Mukherjee & Suppakitpaisarn, 2023; Mohamed et al., 2022).
@@
Beyond the publication of full graphs via randomized response, there is a growing interest in subgraph counting queries under local differential privacy. The challenge was first highlighted by Imola et al. (2021), who proposed two key algorithms for $k$-star and triangle counting. Building on $k$-star counting, Hillebrand et al. (2023) later introduced an unbiased and refined algorithm. For triangle counting, several notable contributions include Imola et al. (2022a), which addressed the communication cost, Liu et al. (2022; 2024b), which delved into scenarios where users have access to a 2-hop graph view, and Eden et al. (2023), which studies the lower bound of the additive error. For larger subgraphs, He et al. (2024) gave a local differentially private algorithm to count butterflies on bipartite graphs, Betzer et al. (2024) gave an algorithm to count the number of walks under the privacy notion, Hillebrand et al. (2025) proposed an algorithm to count odd-length cycles on degeneracy-bounded graphs, and Suppakitpaisarn et al. (2025) devised an algorithm to count any graphlet of size $k$. Moving to the shuffle model, Imola et al. (2022b) gave algorithms for triangle and 4-cycle counting. To conclude, Dhulipala et al. (2022) gave an approximation method for identifying the densest subgraph, rooted in locally adjustable algorithms.
@@
Most of the works mentioned in the previous paragraph use the two-step mechanism (Imola et al., 2021; 2022a; 2022b). In this mechanism, users are required to download the adjacency vectors of all other participants. They then compute graph statistics locally using their genuine adjacency vector and the obfuscated vectors of their peers. Although this technique substantially increases the accuracy for various graph statistics, such as the count of triangles, it comes with the drawback of demanding a vast amount of data download. Each user incurs a communication cost of $\Theta(n^{2})$. Given that the user count, $n$, can reach several billion in real-world scenarios, this download demand is not affordable for most.
@@
Instead of using adjacency vectors, users might consider exchanging adjacency lists. This approach involves each user transmitting and receiving only a list of adjacent node pairs. If we assume that a user can have a limited, constant number of friends, then the bit-download requirement for the original graph becomes $\Theta(n\log n)$. However, given the surge in edge count to $\Theta(n^{2})$ due to the randomized response, the communication cost associated with downloading the obfuscated adjacency list shoots up to $\Theta(n^{2}\log n)$. This means that, post-obfuscation, the adjacency list fails to offer any communication efficiency benefits over the adjacency vector.
@@
There are several techniques proposed to reduce the communication cost for the two-step mechanism such as the asymmetric randomized response (Imola et al., 2022a), the degree-preserving randomized response (Hidano & Murakami, 2024), the degree-preserving exponential mechanism (Adhikari et al., 2020), or technique based on the compressive sensing (Li et al., 2011). However, we strongly believe that the communication cost could be further reduced both in theory and practice.
1.1 Our Contributions
@@
The concept of database compression is not novel within the realm of differential privacy. Numerous mechanisms have been developed to enhance the precision of statistics derived from tabular data through database compression using sampling. One of the most prominent among these is the SmallDB algorithm (Blum et al., 2013). There have also been initiatives that harness the Kronecker graph to refine the accuracy of published graph statistics (Paul et al., 2020).
Figure 1: Overview of our mechanism, GroupRR: the information within the green box represents the data published by users. By the amplification-by-sub-sampling theorem, the randomized response in Step 2.3 has a bit-flip probability reduced by approximately a factor of $s$ relative to the standard randomized response. This helps reducing the error we have in our proposed algorithm.
On the contrary, implementing database compression for the two-step mechanism under edge local differential privacy poses greater challenges. Imagine we compress the adjacency list $a_{i,1},\dots,a_{i,n}]^{t}$ down to $a_{i,c_{1}},\dots,a_{i,c_{m}}]^{t}$, with $m\ll n$, via sampling, and then disseminate this compressed list to all users. When other users utilize this compressed list for local calculations in the second step, they must be able to replicate the sampling outcome. We therefore need a function which can be replicated with small communication cost. This requirement inspired our proposition to utilize linear congruence hashing (Thomson, 1958; Rotenberg, 1960) for the sampling process.
Our proposal is outlined in Figure 1. Using linear congruence hashing, we can evenly partition the node set $\{v_{1},\dots,v_{n}\}$ into $m$ subsets. Let us denote these subsets as $S_{1},\dots,S_{m}$, each having a size of $s$. The index $c_{i}$, corresponding to the only value from the group sent to the server, is derived from $S_{i}$ with a uniform probability of $1/s$. Since all data is transmitted from the users with a probability of $1/s$, we can leverage the theorem on amplification by sub-sampling (Balle et al., 2018) to diminish the bit-flipping probability in the randomized response mechanism by a factor of $1/s$. Consequently, the count of non-zero entries in the bit vector $a_{i,c_{1}},\dots,a_{i,c_{m}}]^{t}$ is roughly $m/s$, which is approximately equal to $n/s^{2}$. By employing a sampling size of $s$, the communication overhead can be reduced by a factor of $s^{2}$.
@@
The main contribution we present in the article is the first purely local differential private mechanism for graph statistics that leverages amplification by sub-sampling. Additionally, we demonstrate the generality of the mechanism and its efficiency as it performs better than state of the art by several orders of magnitudes for several triangle and other subgraph counting tasks.
1.2 Related Works
Hashing functions have been utilized in the domain of differential privacy by Wang et al. (2017) to develop binary local hashing (BLH) and optimal local hashing (OLH) for frequency estimation under local differential privacy. However, the work focuses on tabular information, while our work focuses on graph information. The proposed algorithm is totally different because of the difference in data structure and privacy definition. Additionally, while the previous method use hash functions to compress user data, they do not support amplification by sub-sampling, which is a feature of our proposed method. Thus, our compression rate is larger than the previous work.
A recent study on communication reduction in local differential privacy (Liu et al., 2024a) successfully simulated any randomizer for local differential privacy using only $\mathcal{O}(\varepsilon)$ communication bits. However, similar to BLH and OLH methods, their approach depends on the characteristics of tabular data and traditional local differential privacy, making it inapplicable to our scenario where edge-local differential privacy is employed.
2 Preliminaries
2.1 Hash functions
Hash functions (Knuth, 1997) map keys from a specified key space to a designated hash space. In this paper, unless mentioned differently, the key space is represented by [|0,n1|][|0,n-1|] and the hash space by [|0,m1|][|0,m-1|].
@@ 2.1 Hash functions
We consider a hashing scheme called linear congruence hashing (Thomson, 1958; Rotenberg, 1960) in this paper. The scheme can be defined as $\mathcal{H}=\{h_{\mathsf{a},\mathsf{b}}|\mathsf{a}\in[|1,\mathsf{p}-1|],\mathsf{b}\in[|0,\mathsf{p}-1|]\}$ when $\mathsf{p}$ is a prime number greater than $n$ and $h_{\mathsf{a},\mathsf{b}}(\mathsf{x})=((\mathsf{a}\mathsf{x}+\mathsf{b})\text{ mod }\mathsf{p})\text{ mod }m$. Notice that, for all $\mathsf{a,b}$ and $\mathsf{k}\in[|0,m-1|]$, we have $\left\lfloor\frac{n}{m}\right\rfloor\leq|\{\mathsf{x}:h_{\mathsf{a,b}}(\mathsf{x})=k\}|\leq\left\lceil\frac{n}{m}\right\rceil$. In other words, the number of key in a given bin is either $\left\lfloor\frac{n}{m}\right\rfloor$ or $\left\lceil\frac{n}{m}\right\rceil$.
2.2 Edge Local Differential Privacy
@@
Let the set of users be $\{v_{1},\dots,v_{n}\}$. Each user $v_{i}$ has own adjacency list $a_{i}=[a_{i,1},\dots,a_{i,n}]^{t}\in\{0,1\}^{n}$ where $a_{i,j}=1$ if $\{v_{i},v_{j}\}\in E$ and $a_{i,j}=0$ otherwise. For any $a_{i},a_{i}^{\prime}\in\{0,1\}^{n}$, let $d(a_{i},a_{i}^{\prime})$ be the $\ell_{1}$-distance between the two vectors. To define the privacy notion, we first give the following definition:
Definition 1 (Local differentially private query).
Let $\epsilon>0$. A randomized query $\mathcal{R}$ is said to be $\epsilon$-edge locally differentially private for node $i$ if, for any pair of adjacency lists $a_{i}$ and $a^{\prime}_{i}$ where $|a_{i}-a^{\prime}_{i}|\leq 1$, and for any possible set of outcomes $S$, we have that
$\Pr[\mathcal{R}(a_{i})\in S]\leq e^{\epsilon}\Pr[\mathcal{R}(a_{i}^{\prime})\in S].$
Then, the definition of the edge local differential privacy is as follows:
Definition 2 (Edge local differential privacy Qin et al. (2017)).
An algorithm $\mathcal{A}$ is defined as $\epsilon$-edge locally differentially private if, for any user $i$ and for any possible set of queries $\mathcal{R}_{1},\dots,\mathcal{R}_{\kappa}$ which $\mathcal{A}$ posed to user $i$, where each query $\mathcal{R}_{j}$ is $\epsilon_{j}$-edge locally differentially private (for $1\leq j\leq\kappa$), the condition $\epsilon_{1}+\cdots+\epsilon_{\kappa}\leq\epsilon$ is satisfied.
@@
Theorem 1 (Composition Theorem (Dwork et al., 2010)).
@@
2.3 Basic Mechanisms
Numerous mechanisms have been suggested to meet Definition 2 (Li et al., 2022; Hou et al., 2023). Among them, the edge local Laplacian mechanism (Hillebrand et al., 2023) is an approach to offer privacy when we request each user to give real numbers to the aggregator. The Laplacian mechanism can be defined as in the following definition. It can be shown that the mechanism is $\varepsilon$-edge local differential private.
Definition 3 (Edge Local Laplacian Mechanism (Hillebrand et al., 2023)).
For $f:\{0,1\}^{n}\rightarrow\mathbb{R}$, the global sensitivity of $f$ is $\Delta_{f}=\max\limits_{d(a,a^{\prime})=1}\lVert f(a)-f(a^{\prime})\rVert_{1}$. For $\varepsilon>0$, let $X$ be a random variable drawn from $\textnormal{{Lap}}(\varepsilon/\Delta_{f})$. The edge local Laplacian mechanism for $f$ with privacy budget $\varepsilon$ is the mechanism for which the randomized function $\mathcal{R}$ is defined as $f(a)+X$.
When we request users to send a vector of bits (such as the whole adjacency vector) to the aggregator, the edge local Laplacian mechanism is not the most suitable. Instead, a frequently used method is the randomized response (Warner, 1965; Mangat, 1994; Wang et al., 2016), as detailed in the following definition.
Definition 4 (Randomized Response (Warner, 1965; Mangat, 1994; Wang et al., 2016)).
For $\varepsilon>0$, the randomized response with privacy budget $\varepsilon$ is a mechanism which the randomized function $\mathcal{R}:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$ is a function such that $\mathcal{R}(a_{i,1},\dots,a_{i,n})=(\textnormal{{RR}}(a_{i,1}),\dots,\textnormal{{RR}}(a_{i,n}))$ where
$\mathbb{P}\left(\textnormal{{RR}}(a_{i,j})=1\right)=\begin{cases}\frac{e^{\varepsilon}}{1+e^{\varepsilon}}&\text{if }a_{i,j}=1\\ \frac{1}{1+e^{\varepsilon}}&\text{if }a_{i,j}=0.\end{cases}$ (1)
2.4 Smooth Sensitivity (Nissim et al., 2007)
The traditional Laplacian mechanism adjusts noise based on the global sensitivity, which is determined from the worst possible outcome for a selected function. It does not consider the specific adjacency list. Therefore, in certain situations, the global sensitivity might be much greater than the function’s changes due to slight update in the real adjacency list. Noticing this, Nissim et al. (2007) created a mechanism where the noise relates to the smooth sensitivity, which varies depending on the instance. Before we define smooth sensitivity, we first need to explain local sensitivity and sensitivity at a distance $k$.
Definition 5 (Local Sensitivity).